OpenStudy (anonymous):

divide 15x^5-21x^4+18^2 / 3x^2

OpenStudy (lgbasallote):

\[\Large \frac{15x^5}{3x^2} - \frac{21x^4}{3x^2} + \frac{18x^2}{3x^2}\]

OpenStudy (lgbasallote):

basically...you can divide EACH term of the numerator by the denominator

OpenStudy (anonymous):

after I do that will that be the answer? Or will I have to simplify it more?

OpenStudy (lgbasallote):

that's pretty much it

OpenStudy (anonymous):

Okay well I got 5x^3-7x^2+6x. Does that sound right?

OpenStudy (lgbasallote):

no x in 6x...

OpenStudy (lgbasallote):

x^2 just cancels out

OpenStudy (anonymous):

Oh okay, I questioned myself on that part. Tnks so much :)